A ball on the end of a string is whirled around in a horizontal circle of radius 0.340 m. The plane of the circle is 1.20 m above the ground. The string breaks and the ball lands 2.50 m (horizontally) away from the point on the ground directly beneath the balls location when the string breaks. Find the radial acceleration of the ball during its circular motion. Magnitude m/s Direction
Answer #1

egut 1,2 nead to we haw cre d y 2 0-494Solution) This is a 2D projectile motion problem and so here we used

Yf = Yo + Vyo(t) + (1/2)ay(t^2)

Here Yf is final position

Yo is height or initial position

Vyo is velocity in y direction

ay is acceleration due to gravity

t is time

So finally radial acceleration a = (v^2)/r

Here v = x/t

x is horizontal distance